OpenStudy (anonymous):

if cosx = 3/5 , find exact value of cos2x ?

Parth (parthkohli):

\[\cos(2x) = \cos^2 (x) - \sin^2(x)\]

Parth (parthkohli):

I hope you also know that \(\sin^2(x) = 1 - \cos^2(x)\)

Parth (parthkohli):

So we want\[2\ cos^2(x) + 1\]

geerky42 (geerky42):

Also, \(\cos^2 \theta = \cos \theta \cdot \cos \theta\). And we know that \(\cos \theta= \dfrac{3}{5}\). So you should know what to do next.